# VIRUS-NeRF - Vision, InfraRed and UltraSonic based Neural Radiance Fields

### v1

@@ IV-D2 Discussion
The reduced accuracy of VIRUS-NeRF in the common area compared to the office may be explained by the IRS having a short range of . In the office, 35.7% of all IRS measurements are valid compared to only 10.3% in the common area where the objects are more spread out. However, for safety-relevant tasks, e.g. obstacle detection, the coverage is more important than accuracy and this one remains comparable to LiDAR point clouds in all environments. VIRUS-NeRF predicts more outliers too close leading to false-positive predictions. When visualizing the results as in Fig. 4, the estimation in front of the robot is usually accurate and the hallucinations are either sideways to the trajectory (e.g. in Fig. 3(b)) or further away from the current position of the robot (e.g. in Fig. 3(b)). Lateral to the path only a few measurements are taken causing erroneous predictions. This could be addressed by adding sensors pointing in these directions. Distant hallucinations are most likely caused by the volume rendering, which is biased towards underestimating the depth as explained in chapter III-A3. If the robot moves closer to a particular region, then fewer dispensable samples influence the volume rendering and the false-positive predictions disappear, e.g. compare in Fig. 3(a) and 3(b). In contrast to path planning, hallucinations are not as critical for safety-relevant tasks, e.g. collision avoidance, especially if they vanish when moving closer.
@@ IV-E1 Results
In the ablation study, VIRUS-NeRF is compared to Instant-NGP. Additionally, the contribution of the depth supervision and the improved occupancy grid are studied, and different sensor modalities are analyzed. VIRUS-NeRF is developed and the hyper-parameters are fine-tuned in the office environment and the results of the common area show if the model generalizes well. The mean and inlier percentage are shown in table I. For the common area, the RGB-D camera outperforms all other sensor constellations because its depth images are denser and more accurate than low-cost alternatives (compare Fig. 4(a) and 4(b)). The second-best results are achieved by using VIRUS-NeRF. When removing either the USS or the IRS, the performance drops. The original Instant-NGP implementation is significantly worse in all metrics. When adding the depth losses and still using the occupancy grid of Instant-NGP, the results improve but do not reach the same level as VIRUS-NeRF. The same ablation study is repeated for the office environment: The results are similar, with the exception that omitting pose optimization is less severe or even better in terms of inlier metrics than in the common area, as smaller scenes are less affected by odometry drift. The occupancy grid of Instant-NGP has slightly better mean coverage than VIRUS-NeRF in the office.
@@
/ Sensors | Scene |
Acc. |
Cov. |
Acc. |
Cov. |
Instant-NGP [33] |
C |
0.712 |
1.287 |
0.056 |
0.059 |
CAM |
O | 0.281 | 0.497 | 0.221 | 0.201 |
Instant-NGP |
C |
0.509 |
0.501 |
0.131 |
0.22 |
CAM, USS, IRS |
O | 0.164 |
@@
0.497 | 0.506 |
VIRUS-NeRF |
C |
0.704 |
1.412 |
0.052 |
0.056 |
CAM |
O | 0.277 | 0.476 | 0.231 | 0.215 |
VIRUS-NeRF |
C |
0.49 |
0.568 |
0.146 |
0.146 |
CAM, USS |
O | 0.262 | 0.349 | 0.212 | 0.176 |
VIRUS-NeRF |
C |
0.625 |
0.643 |
0.097 |
0.169 |
CAM, IRS |
O | 0.23 | 0.374 | 0.415 | 0.46 |
VIRUS-NeRF |
C |
@@ IV-E2 Discussion
IV-F Training Speed
All training is done on a Nvidia Titan Xp GPU. The absolute speed of the presented results is expected to be approximately 20% faster if using the Cuda implementation of Instant-NGP instead of the Taichi implementation utilized in this study [39]. Fig. 5(a) shows the average NND over 10 runs for the office environment during training. Using the entire dataset, VIRUS-NeRF converges after around while being significantly faster than Instant-NGP as explained in the following.
IV-F1 Occupancy Grid
On average, VIRUS-NeRF makes 11.64 training steps per second, in contrast to 7.96 when using the grid of Instant-NGP. This is a speed-up of 46% and can be explained by two factors: Instant-NGP samples during the first 256 training steps all grid cells and afterwards a quarter, which is significantly more than 1024 samples used in VIRUS-NeRF. Additionally, the occupancy grid of VIRUS-NeRF relies partially on probabilistic sensor models (i.e. Depth-Update, see section III-B2) instead of inferring with the NeRF network which is computationally less expensive. Accelerating the training process with the occupancy grid of VIRUS-NeRF could be very interesting for any real-time application based on depth supervised Instant-NGP, e.g. NeRF-SLAM [35].

### v2

@@ IV-D2 Discussion
The reduced accuracy of VIRUS-NeRF in the common area compared to the office may be explained by the IRS having a short range of . In the office, 35.7% of all IRS measurements are valid compared to only 10.3% in the common area where the objects are more spread out. However, for safety-relevant tasks, e.g. obstacle detection, the coverage is more important than accuracy and this one remains comparable to LiDAR point clouds in all environments. VIRUS-NeRF predicts more outliers too close leading to false-positive predictions. When visualizing the results as in Fig. 4, the estimation in front of the robot is usually accurate and the hallucinations are either sideways to the trajectory (e.g. in Fig. 4(b)) or further away from the current position of the robot (e.g. in Fig. 4(b)). Lateral to the path only a few measurements are taken causing erroneous predictions. This could be addressed by adding sensors pointing in these directions. Distant hallucinations are most likely caused by the volume rendering, which is biased towards underestimating the depth as explained in chapter III-A3. If the robot moves closer to a particular region, then fewer dispensable samples influence the volume rendering and the false-positive predictions disappear, e.g. compare in Fig. 4(a) and 4(b). In contrast to path planning, hallucinations are not as critical for safety-relevant tasks, e.g. collision avoidance, especially if they vanish when moving closer.
@@ IV-E1 Results
In the ablation study, VIRUS-NeRF is compared to Instant-NGP. Additionally, the contribution of the depth supervision and the improved occupancy grid are studied, and different sensor modalities are analyzed. VIRUS-NeRF is developed and the hyper-parameters are fine-tuned in the office environment and the results of the common area show if the model generalizes well. The mean and inlier percentage are shown in table I. For the common area, the RGB-D camera outperforms all other sensor constellations because its depth images are denser and more accurate than low-cost alternatives (compare Fig. 5(a) and 5(b)). The second-best results are achieved by using VIRUS-NeRF. When removing either the USS or the IRS, the performance drops. The original Instant-NGP implementation is significantly worse in all metrics. When adding the depth losses and still using the occupancy grid of Instant-NGP, the results improve but do not reach the same level as VIRUS-NeRF. The same ablation study is repeated for the office environment: The results are similar, with the exception that omitting pose optimization is less severe or even better in terms of inlier metrics than in the common area, as smaller scenes are less affected by odometry drift. The occupancy grid of Instant-NGP has slightly better mean coverage than VIRUS-NeRF in the office.
@@
/ Sensors | Scene | Acc. | Cov. | Acc. | Cov. |
Instant-NGP [33] | C | 0.712 | 1.287 | 0.056 | 0.059 |
CAM | O | 0.281 | 0.497 | 0.221 | 0.201 |
Instant-NGP | C | 0.509 | 0.501 | 0.131 | 0.22 |
@@
VIRUS-NeRF | C | 0.704 | 1.412 | 0.052 | 0.056 |
CAM | O | 0.277 | 0.476 | 0.231 | 0.215 |
VIRUS-NeRF | C | 0.49 | 0.568 | 0.146 | 0.146 |
CAM, USS | O | 0.262 | 0.349 | 0.212 | 0.176 |
VIRUS-NeRF | C | 0.625 | 0.643 | 0.097 | 0.169 |
CAM, IRS | O | 0.23 | 0.374 | 0.415 | 0.46 |
[…]
VIRUS-NeRF | C | 0.728 | 0.922 | 0.113 | 0.186 |
@@ IV-E2 Discussion
IV-F Training
All training is done on a Nvidia Titan Xp GPU. The algorithm is expected to be approximately 20% faster if using the Cuda implementation of Instant-NGP instead of the Taichi implementation utilized in this study [39].
IV-F1 Speed
Fig. 6(a) shows the average NND over 10 runs for the office environment during training. Using the entire dataset, VIRUS-NeRF converges after around . On average, VIRUS-NeRF makes 11.64 training steps per second, in contrast to 7.96 when using the grid of Instant-NGP. This is a speed-up of 46% and can be explained by two factors: Instant-NGP samples during the first 256 training steps all grid cells and afterwards a quarter, which is significantly more than 1024 samples used in VIRUS-NeRF. Additionally, the occupancy grid of VIRUS-NeRF relies partially on probabilistic sensor models (i.e. Depth-Update, see section III-B2) instead of inferring with the NeRF network which is computationally less expensive. Accelerating the training process with the occupancy grid of VIRUS-NeRF could be very interesting for any real-time application based on depth supervised Instant-NGP, e.g. NeRF-SLAM [35]. To further increase the convergence speed, the same forward pass could be used for the NeRF-Update of the occupancy grid as for the training of the MLP instead of treating these tasks separately. Furthermore, the density MLP could be partially trained without the color MLP leveraging the existence of depth measurements.